OpenStudy (anonymous):

(x+Y) : (x-y)=4:1 so (x2+y2) : (x2-y2)=?

mathslover (mathslover):

are there any options given ?

mathslover (mathslover):

we can solve this further easily but before that i want to know that whether are the options given ? so that i can know that in which terms u want answer

OpenStudy (anonymous):

You are going very difficult @mathslover ..

mathslover (mathslover):

:) that is my way sorry .... let me think of an easy one @waterineyes wait

OpenStudy (anonymous):

Well, what is the profit of your way if the asker does not understand it..

mathslover (mathslover):

sorry have to go now..will answer soon @waterineyes please help him .. @waterineyes i am thinking about this can u wait or just answer

OpenStudy (anonymous):

Ya, I can help you must leave now..

mathslover (mathslover):

i am getting \(\sqrt{236}xy/15\)

OpenStudy (anonymous):

See, we have given; (x + y)/(x - y) = 4/1 So, Cross multiplying we get, 4x - 4y = x + y 3x = 5y Or x = 5y/3 Put this in the required expression: \[(x^2 + y^2)/(x^2 - y^2) = (25y^2/9 + y^2)/(25y^2/9 - y^2) = y^2(25/9 + 1)/y^2(25/9 - 1)\] \[= (25/9 + 1)/(25/9 - 1) = 34/16 = 17/8\] \[(x^2 + y^2)/(x^2 - y^2) = 17/8 = 17:8\]

mathslover (mathslover):

great work @waterineyes

OpenStudy (anonymous):

it can be easily done by using componendo and dividendo property of ratios and proportions.....

OpenStudy (anonymous):

I don't know much about that so I had to go with this method...

OpenStudy (anonymous):

I think it is like: (x + y + x - y)/(x + y - x + y) = (4+1)/(4 - 1) 2x/2y = 5/3 x/y = 5/3 Or Squaring both the sides, \[x^2/y^2 = 25/9\] Using again Componendo And Dividendo, \[(x^2 + y^2)/(x^2 - y^2) = (25 + 9)/(25 - 9) = 34/16 = 17/8\] Yes guys, this is very simple method to solve these type of questions...